# Tony Lindeberg

### v2

@@ Abstract
This paper presents results of combining (i) theoretical analysis regarding connections between the orientation selectivity and the elongation of receptive fields for the affine Gaussian derivative model with (ii) biological measurements of orientation selectivity in the primary visual cortex, to investigate if (iii) the receptive fields can be regarded as spanning a variability in the degree of elongation.
From an in-depth theoretical analysis of idealized models for the receptive fields of simple cells and complex cells in the primary visual cortex, we have established that the directional selectivity becomes more narrow with increasing elongation of the receptive fields. By comparison with previously established biological results, concerning broad vs. sharp orientation tuning of visual neurons in the primary visual cortex, we demonstrate that those underlying theoretical predictions, in combination with these biological results, are consistent with a previously formulated biological hypothesis, stating that the biological receptive field shapes should span the degrees of freedom in affine image transformations, to support affine covariance over the population of receptive fields in the primary visual cortex.
Based on this possible indirect support for the working hypothesis concerning affine covariance, we formulate a set of testable predictions that could be used to, with neurophysiological experiments, judge if the receptive fields in the primary visual cortex of higher mammals could be regarded as spanning a variability over the eccentricity or the elongation of the receptive fields, and, if so, then also characterize if such a variability would, in a structured way, be related to the pinwheel structure in the visual cortex.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 1: Variability in image structures generated by viewing the same surface pattern for different slant angles, that is different angles between the surface normal and the viewing direction. Observe how the resulting perspective effects lead to perspective foreshortening in the vertical direction. (When taking these images, the distance between the object and the observer as well as the amount of zoom for the zoom lens was, however, also adjusted to keep the vertical size of the surface pattern roughly the same in the image domain.)
$\kappa=1$ $\kappa=\sqrt{2}$ $\kappa=2$ $\kappa=2\sqrt{2}$ $\kappa=4$ $\kappa=4\sqrt{2}$
@@
Figure 2: Variability in the elongation of affine Gaussian derivative receptive fields (for image orientation $\varphi=\pi/2$), with the scale parameter ratio $\kappa=\sigma_{2}/\sigma_{1}$ increasing from $1$ to $4\sqrt{2}$ according to a logarithmic distribution, from left to right, with the larger of these scale parameters $\sigma_{2}$ kept constant. (top row) First-order directional derivatives of affine Gaussian kernels according to (1) for $m=1$. (bottom row) Second-order directional derivatives of affine Gaussian kernels according to (1) for $m=2$. (Horizontal axes: image coordinate $x_{1}$. Vertical axes: image coordinate $x_{2}$.)
@@ 1 Introduction
When observing objects and events in our natural environment, the image structures in the visual stimuli will be subject to a substantial variability, caused by the natural image transformations. Specifically, if observing a smooth local surface patch from different viewing directions and viewing distances, this variability can to first order of approximation be approximated by local affine transformations (the derivative of the projective mappings between the different views). Within the 4-D variability of general centered 2-D affine transformations, a 1-D variability in the slant angle of the surface normal relative to the viewing direction does, in terms of covariance properties, correspond to a variability in the elongation of the receptive fields, if we would like the responses to be possible to perfectly match under such variabilities (see Figures 1 and 2 for illustrations).
@@
In (Lindeberg 2021; 2023) we have outlined a framework for how covariance properties with respect to geometric image transformations may constitute a fundamental constraint for the receptive fields in the primary visual cortex of higher mammals, to enable the visual computations to be robust under the variabilities in the image structures generated by the natural image transformations. According to the presented theory, based on axiomatically determined receptive field shapes derived from symmetry properties that reflect structural properties of the environment, in combination with additional constraints to guarantee consistency between image representations over multiple spatial and temporal scales, the population of receptive fields in the primary visual cortex ought to, according to this theory, obey covariance properties with respect to spatial affine transformations and Galilean transformations.
@@
While overall qualitative comparisons between predictions from this principled theory have been successfully made to neurophysiological recordings of receptive fields of simple cells by DeAngelis et al. (1995; 2004), Conway and Livingstone (2006) and Johnson et al. (2008), the publicly available data regarding full receptive field recordings are quite limited, why further experimental evidence would be needed to firmly either reject or support the stated hypotheses about affine covariance and Galilean covariance.
@@
In the lack of such neurophysiological data regarding full receptive field recordings, one could, however, aim to instead obtain indirect cues regarding a possible variability in the degree of elongation of the receptive fields, by making use of the recordings of the orientation selectivity of visual neurons by Nauhaus et al. (2008), which show a substantial variability regarding broad vs. sharp tuning of the receptive fields in the primary visual cortex.
@@
In a companion paper (Lindeberg 2024), we have established a strong direct link between the orientation selectivity and the elongation of the receptive fields according to the idealized generalized Gaussian derivative model for visual receptive fields (as will be summarized in Section 2). If we would assume that that model would constitute a sufficiently valid model for the population of simple and complex cells in the primary visual cortex, then we could by logical inference infer possible indirect support for the working hypothesis, in that the variability in the orientation selectivity of the receptive fields would correspond to a variability in the degree of elongation of the receptive fields.
@@ 1.1 The hypothesis about affine covariant receptive fields
If we assume that the visual system should implement affine covariant receptive fields, to make it possible for the visual system to better estimate shape properties of surfaces in the world, such as local surface orientation, then the property of affine covariance would make it possible to compute better estimates of local surface orientation, compared to a visual system that does not implement affine covariance, or a sufficiently good approximation thereof.
@@
Fundamentally, we cannot expect that visual perception should be able to implement full affine invariance. For example, from the well-known experience, that it is much harder to read a text upside-down, it is clear that visual perception cannot be regarded as invariant to spatial rotations in the image domain. From the expansion of the orientations of visual receptive fields according to the pinwheel structure of higher mammals, we can, however, regard the population of receptive fields as supporting local rotational covariance. If we look at a slanted surface in the world, we can get a robust perception of its surface texture under substantial variations of the slant angle, thus showing a robustness of visual perception under stretchings of visual image patterns that correspond to non-uniform scaling transformations (the perspective effects on a slanted surface patch can to first-order of approximation be modelled as a stretching of the image pattern along the tilt direction in image space, complemented with a uniform scaling transformation). If the visual receptive fields would span a variability under such spatial stretching transformations, then such a variability would precisely correspond to a variability in the anisotropy, or the degree of elongation, of the receptive fields.
1.2 Variability over the elongation of receptive fields
@@
The overall theme of this article is thus to, based on a theoretical analysis of a relationship between the orientation selectivity and the degree of anisotropy or degree of elongation of receptive fields in the primary visual cortex, in combination with existing neurophysiological results concerning variabilities in the orientation selectivity of the receptive fields in relation to the pinwheel structure of higher mammals, address the question of whether the receptive fields in the primary visual cortex could be regarding as spanning a variability over the elongation of the receptive fields, to, in turn, support covariant image measurements under such variabilities, or, at least, sufficiently good approximations thereof.
@@
If we, in particular, could assume that the spatial components of the receptive fields could be well modelled by affine Gaussian derivatives, then the property of affine covariance implies that there should be visual receptive fields for different degrees of anisotropy present in the visual system. If we further combine the theoretical results used in this paper, which will state that the degree of orientation selectivity is strongly dependent on the anisotropy or the elongation of the receptive fields, with the biological results established by Nauhaus et al. (2008), which show that the degree of orientation selectivity for neurons in the primary visual cortex varies both strongly and in relationship with the position on the cortical surface in relation to the pinwheel structure, then these results together are fully consistent with the hypothesis that the visual receptive fields in the primary visual cortex should span a variability in their anisotropy, thus consistent with the hypothesis that the receptive fields in the pinwheel structure should span at least one more degree of freedom in the affine group, beyond mere rotations (as already established in previous neurophysiological measurements regarding the pinwheel structure of the oriented receptive fields in the primary visual cortex of higher mammals).
@@
We will also, more generally, use predictions from the presented theoretical analysis to formulate a set of explicit testable hypotheses, that could be either verified or rejected in future neurophysiological experiments. Additionally, we will formulate a set of quantitive measurements to be made, to characterize a possible variability in the anisotropy or elongation of receptive fields in the primary visual cortex, with special emphasis on the relationships between a possibly predicted variability in receptive field elongation and the pinwheel structure in the primary visual cortex of higher mammals.
@@
2 Connections between the orientation selectivity and the elongation of receptive fields for the generalized Gaussian derivative model for visual receptive fields
@@
For modelling the receptive fields in the primary visual cortex, we will make use of the generalized Gaussian derivative model for receptive fields.
2.1 Idealized models for simple cells
@@
The purely spatial component of the receptive fields for the simple cells modelled as (Lindeberg 2021 Equation (23); see Figure 7 in that reference for illustrations)
@@
  • $\partial_{\bar{t}}^{n}=(\partial_{t}+v_{1}\,\partial_{x_{1}}+v_{2}\,\partial_{x_{2}})^{n}$ is an $n$:th-order velocity-adapted temporal derivative operator,
@@
This model builds upon the regular Gaussian derivative model for purely spatial receptive fields proposed by Koenderink and van Doorn (1984; 1987; 1992) and previously used for modelling biological fields by Young and his co-workers 1987; 2001; 2001, however, here also generalized to affine covariance according to Lindeberg (2013a; 2021).
2.2 Idealized models for complex cells
To model complex cells with a purely spatial dependency, we will use a quasi quadrature measure of the form (Lindeberg 2020 Equation (39))

### v7

@@ Abstract
This paper presents the results of combining (i) theoretical analysis regarding connections between the orientation selectivity and the elongation of receptive fields for the affine Gaussian derivative model with (ii) biological measurements of orientation selectivity in the primary visual cortex to investigate if (iii) the receptive fields can be regarded as spanning a variability in the degree of elongation.
From an in-depth theoretical analysis of idealized models for the receptive fields of simple and complex cells in the primary visual cortex, we established that the orientation selectivity becomes more narrow with increasing elongation of the receptive fields. Combined with previously established biological results, concerning broad vs. sharp orientation tuning of visual neurons in the primary visual cortex, as well as previous experimental results concerning distributions of the resultant of the orientation selectivity curves for simple and complex cells, we show that these results are consistent with the receptive fields spanning a variability over the degree of elongation of the receptive fields. We also show that our principled theoretical model for visual receptive fields leads to qualitatively similar types of deviations from a uniform histogram of the resultant descriptor of the orientation selectivity curves for simple cells, as can be observed in the results from biological experiments.
To firmly determine if the underlying working hypothesis, regarding the receptive fields spanning a variability in the degree of elongation, would truly hold for the receptive fields in the primary visual cortex of higher mammals, we formulate a set of testable predictions, that can be used for investigate this property experimentally, and, if applicable, then also characterize if such a variability would, in a structured way, be related to the pinwheel structure in the visual cortex.
@@
Refer to caption Refer to caption Refer to caption
Figure 1: Variabilities in image structures generated by viewing the same surface patterns from different viewing directions. Observe how the resulting perspective transformations lead to strong foreshortening effects, in that the image structures in one direction in the 2-D image space are compressed more than the image structures in the orthogonal direction. Here, where the viewpoint of the observer is moved horizontally in the world, the foreshortening effect is mainly along the horizontal direction, although complemented also with small rotations for non-central image points, because of using a planar image plane as opposed to a spherical retina. To first order of approximation of the projective mappings between pairwise views, these resulting image deformations can be modelled in terms of local affine transformations.
$\kappa=1$ $\kappa=2\sqrt{2}$ $\kappa=2$ $\kappa=2\sqrt{2}$ $\kappa=4$ $\kappa=4\sqrt{2}$
@@
Figure 2: Variability in the elongation of affine Gaussian derivative receptive fields (for image orientation $\varphi=0$), with the scale parameter ratio $\kappa=\sigma_{2}/\sigma_{1}$ increasing from $1$ to $4\sqrt{2}$ according to a logarithmic distribution, from left to right, with the vertical scale parameter kept constant $\sigma_{2}=4$ and with the horizontal scale parameter being the smaller $\sigma_{1}\leq\sigma_{2}$. (first row) First-order directional derivatives of affine Gaussian kernels according to (5) for $m=1$. (second row) Second-order directional derivatives of affine Gaussian kernels according to (5) for $m=2$. (third row) Second-order directional derivatives of affine Gaussian kernels according to (5) for $m=3$. (fourth row) Second-order directional derivatives of affine Gaussian kernels according to (5) for $m=4$. When the image structures are subject to foreshortening transformations, because of varying slant angles between the local surface normals and the viewing direction, the shapes of the spatial components of the receptive fields need to be adapted accordingly, to achieve affine covariance in the sense that the receptive field responses should be possibly to be appropriately matched under different viewing conditions. (Horizontal axes: image coordinate $x_{1}\in[-16,16]$. Vertical axes: image coordinate $x_{2}\in[-16,16]$.)
@@ 1 Introduction
When observing objects and events in our natural environment, the image structures in the visual stimuli will be subject to substantial variabilities caused by the natural image transformations. Specifically, if observing a smooth local surface patch from different viewing directions and viewing distances, this variability can, to first order of approximation, be approximated by local affine transformations (the derivative of the projective mappings between the different views). Within the 4-D variability of general centered 2-D affine transformations, a 1-D variability in the slant angle of the surface normal relative to the viewing direction does, in terms of covariance properties, correspond to a variability in the elongation of the receptive fields, if we would like the responses to be possible to perfectly match under such variabilities (see Figures 1 and 2 for illustrations and Section 2.1 for a complementary theoretical background).
@@
In Lindeberg (2021; 2023; 2024b), we outlined a framework for how covariance properties with respect to geometric image transformations may constitute a fundamental constraint for the receptive fields in the primary visual cortex of higher mammals, to enable the visual computations to be robust under the variabilities in the image structures generated by the natural image transformations. According to the presented theory, based on axiomatically determined receptive field shapes derived from symmetry properties that reflect structural properties of the environment, in combination with additional constraints to guarantee consistency between image representations over multiple spatial and temporal scales, the population of receptive fields in the primary visual cortex ought to, according to this theory, obey covariance properties with respect to spatial affine transformations and Galilean transformations.
@@
While overall qualitative comparisons between predictions from this principled theory have been successfully made to neurophysiological recordings of receptive fields of simple cells by DeAngelis et al. (1995; 2004), Conway and Livingstone (2006) and Johnson et al. (2008), publicly available data regarding full receptive field recordings are quite limited. Thus, further experimental evidence would be needed to firmly either reject or support the stated hypotheses about affine covariance and Galilean covariance.
@@
With the lack of such neurophysiological data regarding full receptive field recordings, one could, however, aim to instead obtain indirect cues regarding a possible variability in the degree of elongation of the receptive fields, by making use of the recordings of the orientation selectivity of visual neurons by Nauhaus et al. (2008), which show a substantial variability regarding broad vs. sharp tuning of the receptive fields in the primary visual cortex, as well as by Goris et al. (2015), who report comparably uniform distributions of the degree of orientation selectivity of simple and complex cells, in terms of histograms of the resultant of the orientation selectivity curves.
@@
In a companion paper (Lindeberg 2024a), we established a strong direct link between the orientation selectivity and the elongation of the receptive fields according to the idealized generalized Gaussian derivative model for visual receptive fields (as will be summarized in Section 2.2). If we would assume that the generalized Gaussian derivative model would constitute a sufficiently valid model for the population of simple and complex cells in the primary visual cortex, then we could logically infer possible indirect support for the working hypothesis, in that the observed variability in the orientation selectivity of the receptive fields would correspond to a variability in the degree of elongation of the receptive fields.
@@ 1.1 The hypothesis about affine covariant receptive fields
If we assume that the visual system should implement affine covariant receptive fields (Lindeberg 2023 Section 3.2), then the property of affine covariance would make it possible to compute better estimates of local surface orientation, compared to a visual system that does not implement affine covariance, or a sufficiently good approximation thereof.
@@
Fundamentally, we cannot expect the visual perception system to implement full affine invariance. For example, from the well-known experience, that it is much harder to read text upside-down, it is clear that the visual perception system cannot be regarded as invariant to spatial rotations in the image domain. However, from the expansion of the orientations of visual receptive fields according to the pinwheel structure of higher mammals, we can regard the population of receptive fields as supporting local rotational covariance.
When we look at a slanted surface in the world, we can get a robust and stable perception of its surface texture under substantial variations of the slant angle. This robustness of the visual perception system under stretchings of image patterns that correspond to non-uniform scaling transformations (the perspective effects on a slanted surface patch can, to first-order of approximation, be modelled as a stretching of the image pattern along the tilt direction in image space, complemented with a uniform scaling transformation). If the visual receptive fields would span a variability under such spatial stretching transformations, then such a variability would precisely correspond to a variability in the anisotropy, or the degree of elongation, of the receptive fields.
1.2 Variability over the degree of elongation of receptive fields
@@
The overall theme of this study is to, based on a theoretical analysis of a relationship between the orientation selectivity and the degree of anisotropy or degree of elongation of the receptive fields in the primary visual cortex, in combination with existing neurophysiological results concerning variabilities in the orientation selectivity of the receptive fields in relation to the pinwheel structure of higher mammals, address the question of whether the receptive fields in the primary visual cortex could be regarded as spanning a variability over the elongation of the receptive fields, to support covariant image measurements under such variabilities, or, at least, sufficiently good approximations thereof.
@@
In particular, if we could assume that the spatial components of the receptive fields could be well modelled by affine Gaussian derivatives, then the property of affine covariance implies that there should be visual receptive fields for different degrees of anisotropy present in the visual system. If we further combine the theoretical results used in this paper, which state that the degree of orientation selectivity is strongly dependent on the anisotropy or elongation of the receptive fields, with the biological results established by Nauhaus et al. (2008), which show that the degree of orientation selectivity for neurons in the primary visual cortex varies both strongly and in relationship with the position on the cortical surface in relation to the pinwheel structure. Then, these results together are fully consistent with the hypothesis that the visual receptive fields in the primary visual cortex should span a variability in their anisotropy, thus consistent with the hypothesis that the receptive fields in the pinwheel structure should span at least one more degree of freedom in the affine group, beyond mere rotations (as already established in previous neurophysiological measurements regarding the pinwheel structure of the oriented receptive fields in the primary visual cortex of higher mammals).
@@
We will also, more generally, use predictions from the presented theoretical analysis to formulate a set of explicit, testable hypotheses, that could be either verified or rejected in future neurophysiological experiments. Additionally, we will formulate a set of quantitive measurements to be made, to characterize a possible variability in the anisotropy or elongation of receptive fields in the primary visual cortex, with special emphasis on the relationships between a possibly predicted variability in receptive field elongation and the pinwheel structure in the primary visual cortex of higher mammals.
1.3 Contributions and novelty
In summary, the purposes of this paper are twofold:
  • to, in the current absence of biological measurements, about a possible variability of the degree of elongation of receptive field shapes in the primary visual cortex, provide potential indirect support for such a hypothesis, based on a combination of previously established variability in the degree of the orientation selectivity of biological receptive fields with a model-based connection between the degree of elongation of the receptive fields and their orientation selectivity, and to express possible indirect support for that hypothesis, and
  • to formulate a set of theoretically motivated and experimentally testable predictions and quantitative measurements, that could be used by experimentalists for ultimately judging whether the formulated hypothesis about a variability over the degree of elongation of the receptive fields would hold in the primary visual cortex of higher mammals, to, in turn, if that hypothesis would hold, provide potential further support for the more general hypothesis, that the receptive fields in the primary visual cortex of higher mammals may be covariant under the larger group of spatial affine transformations.
The main novel contributions of this paper are specifically:
  • A qualitative explanation of the experimental results by Nauhaus et al. (2008) as the results of a systematic variability in the degree of elongation of the receptive fields.
  • Biological implications of that result, in that the pinwheel structure of higher mammals should, beyond an expansion over image orientations, also comprise an expansion over the degree of elongation of the receptive fields, from the center of each pinwheel to its periphery.
  • Theoretical connections to, as well as potential partial support for, the hypothesis that the receptive fields in the primary visual cortex may be covariant under spatial affine transformations.
  • The quantitative explanation of the receptive field histograms obtained experimentally by Goris et al. (2015), as the results of combining the receptive fields of simple cells corresponding to different orders of spatial differentiation.
  • Biological implications of that result, that the experimentally recorded receptive field histograms are better explained by receptive fields in terms of spatial derivatives up to order 4 than in terms of spatial derivatives up to order 2.
  • The overall theoretical contributions in the paper, of explaining properties of the primary visual cortex, based on properties of a theoretically principled model for the receptive fields.
A further underlying motivation of this work is to lay out a conceptual foundation, by which theoreticians and experimentalists could join efforts, to establish to what extent the distributions of the shapes of the biological receptive fields would be compatible with an explanation from the fundamental constraint, that the family of receptive fields should be able to handle variabilities in the image data caused by geometric image transformations.
2 Methods
In this section, we: (i) give a theoretical background regarding the notion of affine covariant visual receptive fields, which constitutes the conceptual background for the hypotheses studied in this work; (ii) describe how the orientation selectivity of receptive fields is related to the degree of elongation of the receptive fields, based on an in-depth theoretical analysis of visual receptive fields according to the generalized Gaussian derivative model; and (iii) relate the computational modelling approaches taken and the contributions presented in this work to previous work in the field.
2.1 Affine covariant visual receptive fields
Let us represent spatial coordinates by $x=(x_{1},x_{2})^{T}$ and centered affine spatial transformations in the 2-D image domain as
$x^{\prime}=A\,x,$ (1)
where $A$ represents any non-singular $2\times 2$ matrix.
Then, an affine transformed image $f^{\prime}(x^{\prime})$ of an original image $f(x)$ is defined according to
$f^{\prime}(x^{\prime})=f(x).$ (2)
With the affine transformation operator ${\cal T}_{A}$, we can write this relationship as
$f^{\prime}={\cal T}_{A}\,f.$ (3)
The property of affine covariance then means that the results of either:
  • applying an affine transformation $x^{\prime}=A\,x$ to an image $f(x)$ and then applying a receptive field ${\cal R}^{\prime}$ to the affine transformed image $f^{\prime}(x^{\prime})$, or
  • applying a related receptive field ${\cal R}$ to the original image $f(x)$ and then applying an affine transformation ${\cal T}_{A}$ to that output,
will lead to the same result, such that
${\cal R}^{\prime}\,{\cal T}_{A}\,f={\cal T}_{A}\,{\cal R}\,f,$ (4)
where the affine covariant property of the receptive field family means that for every receptive field ${\cal R}$ in the receptive field family, there exists a possibly transformed receptive field ${\cal R}^{\prime}$ within the same family, specifically determined according to the actual value of the affine transformation matrix $A$, such that the above relationship is guaranteed to hold, for some transformed receptive field ${\cal R}^{\prime}$ as a function of the original receptive field ${\cal R}$ and the affine transformation matrix $A$.
The property of affine covariance thus means that the family of receptive fields is well-behaved with regard to spatial affine transformations, in the sense that affine transformations commute with the operation of computing outputs from the family of receptive fields.
In Lindeberg (2021; 2023; 2024b), it is argued that such affine covariant properties constitute an essential property of spatial receptive fields, as well as for the spatial components in joint spatio-temporal receptive fields. Specifically, the receptive fields, according to the generalized Gaussian derivative model for visual receptive fields, to be used below, obey such affine covariant properties.
The property of the degree of elongation of the receptive fields, to be studied in detail in this work, spans a 1-D variability within the full 4-D variability of general affine transformations, thus constituting one of the degrees of freedom in the variability of transformed receptive field shapes of ${\cal R}^{\prime}$, that will be generated by subjecting an original receptive field ${\cal R}$ to the 4-D variability of general affine transformation matrices $A$.
@@ 2.1 Affine covariant visual receptive fields
2.2 Connections between the orientation selectivity and the degree of elongation of the receptive fields for the generalized Gaussian derivative model for visual receptive fields
@@
For modelling the receptive fields in the primary visual cortex, we will use the generalized Gaussian derivative model for receptive fields (Lindeberg 2021).
2.2.1 Idealized models for simple cells
@@
We will model the purely spatial component of the receptive fields for simple cells as (Lindeberg 2021 Equation (23); see Figure 7 in that reference for illustrations)
@@
  • $\partial_{\bar{t}}^{n}=(\partial_{t}+v_{1}\,\partial_{x_{1}}+v_{2}\,\partial_{x_{2}})^{n}$ is an $n$:th-order velocity-adapted temporal derivative operator, and
@@
This model builds upon the regular Gaussian derivative model for purely spatial receptive fields proposed by Koenderink and van Doorn (1984; 1987; 1992) and previously used for modelling biological fields by Young and his co-workers (1987; 2001; 2001). Here, that regular Gaussian derivative model is additionally generalized to affine covariance, according to Lindeberg (2013a; 2021).
2.2.2 Idealized models for complex cells
To model complex cells with a purely spatial dependency, we will use a quasi-quadrature measure of the form (Lindeberg 2020 Equation (39))